8.46 A random sample of 10 miniature Tootsie Rolls was taken from a bag. Each piece was weighed on
a very accurate scale.
The results in grams were 3.087 3.131 3.241 3.241 3.270 3.353 3.400 3.411 3.437 3.477
(a) Construct a 90 percent confidence interval for the true mean weight. (b) What sample size
would be necessary to estimate the true weight with an error of ± 0.03 grams with 90 percent confidence?
(c) Discuss the factors which might cause variation in the weight of Tootsie Rolls during
manufacture. (Data are from a project by MBA student Henry Scussel.)

Answer by stanbon(75887)

8.46 A random sample of 10 miniature Tootsie Rolls was taken from a bag. Each piece was weighed on
a very accurate scale.
The results in grams were 3.087 3.131 3.241 3.241 3.270 3.353 3.400 3.411 3.437 3.477
(a) Construct a 90 percent confidence interval for the true mean weight.
sample mean = 3.3048 ; sample std = 0.13199
standard error: 1.645*0.13199/Sqrt(10) = 0.06866
---
90% CI: 3.3048-0.06866 < u < 3.3048+0.06866
90% CI:3.23614 < u < 3.37346
====================================================

(b) What sample size would be necessary to estimate the true weight with an error of ± 0.03 grams with 90 percent confidence?
---
n = [1.645*0.13199/0.03]^2 = 7.24^2 = 52.38
Round up to get n = 53
====================================================

(c) Discuss the factors which might cause variation in the weight of Tootsie Rolls during manufacture. (Data are from a project by MBA student Henry Scussel.)
I'll leave that to you.
===============================
Cheers,
Stan H.